# Difficulty-Aware Agentic Orchestration for Query-Specific Multi-Agent Workflows

### v4

@@ Abstract.
Large Language Model (LLM)-based agentic systems have shown strong capabilities across various tasks. However, existing multi-agent frameworks often rely on static or task-level workflows, which either over-process simple queries or underperform on complex ones, while also neglecting the efficiency-performance trade-offs across heterogeneous LLMs. To address these limitations, we propose Difficulty-Aware Agentic Orchestration (DAAO), which can dynamically generate query-specific multi-agent workflows guided by predicted query difficulty. DAAO comprises three interdependent modules: a variational autoencoder (VAE) for difficulty estimation, a modular operator allocator, and a cost- and performance-aware LLM router. A self-adjusting policy updates difficulty estimates based on workflow success, enabling simpler workflows for easy queries and more complex strategies for harder ones. Experiments on six benchmarks demonstrate that DAAO surpasses prior multi-agent systems in both accuracy and inference efficiency, validating its effectiveness for adaptive, difficulty-aware reasoning.
Multi-agent system, Difficulty-Aware, Adaptive workflows
$c$s: Computing methodologies Cooperation and coordinationccs: Computing methodologies Multi-agent systems
@@ 1. Introduction
Large Language Model (LLM)-based agents (Richards and et al., 2023; Nakajima, 2023; Reworkd, 2023) have exhibited remarkable capabilities across a wide spectrum of tasks, including question answering (Zhu et al., 2024), data analysis (Hong et al., 2024; Li et al., 2024), decision-making (Song et al., 2023), code generation (Shinn et al., 2023) and web navigation (Deng et al., 2024). Building upon the success of single agents, recent advancements reveal that organizing multiple LLM-based agents into structured agentic workflows can significantly enhance task performance. In such workflows, agents can interact either cooperatively (Zhuge et al., 2024) or competitively (Zhao et al., 2023) depending on the task context. These multi-agent systems can overcome the cognitive and functional limitations of individual models (Du et al., 2023; Liang et al., 2023; Wang et al., 2023a; Jiang et al., 2023; Wu et al., 2023; Zhang et al., 2024b), thereby exhibiting collective intelligence similar to human collaboration in a society of agents.
@@ 4.1. Experiment Setup
We evaluate DAAO on six public benchmarks covering three domains: (1) math reasoning, GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b); (2) code generation, HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021)); tool use, GAIA (Mialon et al., 2023). Additionally, we include MMLU (Hendrycks et al., 2021a), a benchmark covering 57 academic subjects, to assess general knowledge and multitask language understanding. For the MATH benchmark, we follow (Hong et al., 2024) in selecting a harder subset (617 problems). The dataset metric are in Appendix C.
@@ 4.4. Inductive Ability Analysis
Dataset
HumanEval MATH
Metric
Pass@1
Cost
($10^{-3}\mathdollar$)
Accuracy
Cost
($10^{-3}\mathdollar$)
Vanilla
$94.65$ $1.10$ $55.37$ $0.55$
@@ 5. Conclusion
In this work, we presented DAAO, a difficulty-aware agentic orchestration framework that dynamically adapts reasoning workflows to the complexity and domain characteristics of each query. By combining query-level difficulty estimation, modular operator allocation, and heterogeneous LLM routing, DAAO constructs flexible and cost-efficient agentic workflows. Our approach moves beyond static, one-size-fits-all designs by leveraging the complementary strengths of diverse LLMs and adapting workflow depth on a per-query basis. Extensive experiments across six benchmarks demonstrate that DAAO consistently outperforms existing multi-agent and LLM routing systems in both accuracy and efficiency, achieving up to 11.21% higher accuracy while reducing inference cost by up to 36%. These results validate the importance of difficulty-guided, modular orchestration in building scalable and performant LLM-based agents. Future work includes extending DAAO to handle multi-modal queries and incorporating real-time feedback for online adaptation.

### v5

@@ Abstract.
Large Language Model (LLM)-based agentic systems have shown strong capabilities across various tasks. However, existing multi-agent frameworks often rely on static or task-level workflows, which either over-process simple queries or underperform on complex ones, while also neglecting the efficiency-performance trade-offs across heterogeneous LLMs. To address these limitations, we propose Difficulty-Aware Agentic Orchestration (DAAO), which can dynamically generate query-specific multi-agent workflows guided by predicted query difficulty. DAAO comprises three interdependent modules: a variational autoencoder (VAE) for difficulty estimation, a modular operator allocator, and a cost- and performance-aware LLM router. A self-adjusting policy updates difficulty estimates based on workflow success, enabling simpler workflows for easy queries and more complex strategies for harder ones. Experiments on six benchmarks demonstrate that DAAO surpasses prior multi-agent systems in both accuracy and inference efficiency, validating its effectiveness for adaptive, difficulty-aware reasoning.Our code is open-sourced at https://github.com/AutoAgents-ai/DAAO
Multi-agent system; Difficulty-Aware; Adaptive workflows
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emirates.booktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesisbn: 979-8-4007-2307-0/2026/04doi: 10.1145/3774904.3792240ccs: Computing methodologies Cooperation and coordinationccs: Computing methodologies Multi-agent systems
@@ 1. Introduction
Large Language Model (LLM)-based agents (Richards and et al., 2023; Nakajima, 2023; Reworkd, 2023; Lou et al., 2025) have exhibited remarkable capabilities across a wide spectrum of tasks, including question answering (Zhu et al., 2024), data analysis (Hong et al., 2024; Li et al., 2024; Shen and Zhang, 2025), decision-making (Song et al., 2023; Liu et al., 2026; Sun et al., 2025), code generation (Shinn et al., 2023) and web navigation (Deng et al., 2024). Building upon the success of single agents, recent advancements reveal that organizing multiple LLM-based agents into structured agentic workflows can significantly enhance task performance. In such workflows, agents can interact either cooperatively (Zhuge et al., 2024) or competitively (Zhao et al., 2023) depending on the task context. These multi-agent systems can overcome the cognitive and functional limitations of individual models (Du et al., 2023; Liang et al., 2023; Wang et al., 2023a; Jiang et al., 2023; Wu et al., 2023; Zhang et al., 2024b), thereby exhibiting collective intelligence similar to human collaboration in a society of agents.
@@ 4.1. Experiment Setup
We evaluate DAAO on six public benchmarks covering three domains: (1) math reasoning, GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b); (2) code generation, HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021)); GAIA (Mialon et al., 2023). Additionally, we include MMLU (Hendrycks et al., 2021a), a benchmark covering 57 academic subjects, to assess general knowledge and multitask language understanding. For the MATH benchmark, we follow (Hong et al., 2024) in selecting a harder subset (617 problems). The dataset metric are in Appendix C.
@@ 4.4. Inductive Ability Analysis
Dataset HumanEval MATH
Metric Pass@1 (%) Cost ($10^{-3}\mathdollar$) Accuracy (%) Cost ($10^{-3}\mathdollar$)
Vanilla $94.65$ $1.10$ $55.37$ $0.55$
@@ 5. Conclusion
In this work, we propose DAAO, a difficulty-aware agentic orchestration framework that dynamically adapts reasoning workflows to query complexity and domain. By estimating query difficulty, allocating modular operators, and routing to heterogeneous LLMs, DAAO builds flexible and cost-efficient workflows. It moves beyond one-size-fits-all designs by leveraging complementary LLM strengths and adapting workflow depth per query. Experiments across six benchmarks show that DAAO outperforms existing multi-agent and routing systems, achieving up to 11.21% higher accuracy with up to 36% lower cost. These results highlight the value of difficulty-guided modular orchestration for scalable LLM agents. Future work includes extending DAAO to multi-modal queries and incorporating real-time feedback.
6. Acknowledgement
This work was supported by the National Natural Science Foundation of China under Grant 52308250 and Autoagents.AI.